John Fleming, chief administrator for Valley View Hospital, is concerned about the costs for tests in the hospital’s lab. Charges for lab tests are consistently higher at Valley View than at other hospitals and have resulted in many complaints. Also, because of strict regulations on amounts reimbursed for lab tests, payments received from insurance companies and governmental units have not been high enough to cover lab costs. Mr. Fleming has asked you to evaluate costs in the hospital’s lab for the past month. The following information is available: Two types of tests are performed in the lab—blood tests and smears. During the past month, 1,300 blood tests and 4,000 smears were performed in the lab. Small glass plates are used in both types of tests. During the past month, the hospital purchased 20,000 plates at a cost of $79,800. 3,000 of these plates were unused at the end of the month; no plates were on hand at the beginning of the month. During the past month, 3,000 hours of labor time were recorded in the lab at a cost of $33,900. The lab’s variable overhead cost last month totaled $25,800. Valley View Hospital has never used standard costs. By searching industry literature, however, you have determined the following nationwide averages for hospital labs: Plates: Three plates are required per lab test. These plates cost $4.20 each and are disposed of after the test is completed. Labor: Each blood test should require 0.8 hours to complete, and each smear should require 0.40 hours to complete. The average cost of this lab time is $12.20 per hour. Overhead: Overhead cost is based on direct labor-hours. The average rate for variable overhead is $8.10 per hour. Required: 1. Compute a materials price variance for the plates purchased last month and a materials quantity variance for the plates used last month. 2. For labor cost in the lab: a. Compute a labor rate variance and a labor efficiency variance. b. In most hospitals, one-half of the workers in the lab are senior technicians and one-half are assistants. In an effort to reduce costs, Valley View Hospital employs only one-fourth senior technicians and three-fourths assistants. Would you recommend that this policy be continued? 3-a. Compute the variable overhead rate and efficiency variances. 3-b. Is there any relation between the variable overhead efficiency variance and the labor efficiency variance?
Answer #1

Solution 1:

Standard quantity of material for actual test = (1300+4000) * 3 = 15900 plates

Actual quantity of material purchased = 20000 plates

Actual quantity of material used = 20000 - 3000 = 17000 plates

Standard price of material = $4.20

Actual price of material = $79,800 / 20000 = $3.99

Material price variance = (SP - AP) * AQ purchased = ($4.20 - $3.99) * 20000 = $4,200 F

Material quantity variance = (SQ - AQ) * SR = (15900 - 17000) * $4.20 = $4,620 U

Solution 2a:

Standard hours of direct labor = 1300*0.8 + 4000*0.4 = 2640 hours

Standard rate of direct labor = $12.20 per hour

Actual hours of direct labor = 3000

Actual rate of direct labor = $33,900 / 3000 = $11.30 per hour

Direct labor rate variance = (SR - AR) * AH = ($12.20 - $11.30) * 3000 = $2,700 F

Direct labor efficiency variance = (SH - AH) * SR = (2640 - 3000) * $12.20 = $4,392 U

Solution 2b:

As a standard, one-half of the workers in the lab are senior technicians and one-half are assistants. However valley view hospital employs only one-fourth senior technicians and three-fourths assistants. This is resulting in favorable direct labor rate variance as direct labor rate per hour is $11.30 from standard rate of $12.20 per hour. However as hospital employs more assistant, it is resulting in to extra time taken by assistant from standard time because they in experienced resulting in unfavorable labor efficiency variance. Unfavorable labor efficiency variance is higher than favorable direct labor rate variance, therefore this policy should be discontinued.

Solution 3a:

Standard hours of direct labor = 1300*0.8 + 4000*0.4 = 2640 hours

Standard rate of variable overhead= $8.10 per hour

Actual hours of direct labor = 3000

Actual rate of variable overhead = $25,800 / 3000 = $8.60 per hour

Variable overhead rate variance = (SR - AR) * AH = ($8.10 - $8.60) * 3000 = $1,500 U

Variable overhead efficiency variance = (SH - AH) * SR = (2640 - 3000) * $8.10 = $2,916 U

Solution 3b:

Yes, there is relation between variable overhead efficiency variance and the labor efficiency variance. Variable overhead is based on direct labor hours, therefore if direct labor efficiency variance improves, variable overhead efficiency variance also improve. Both are positively corelated.